Kingsport Containers Company makes a single product that is subject to wide seasonal variations in demand. The company uses a job-order costing system and computes plantwide predetermined overhead rates on a quarterly basis using the number of units to be produced as the allocation base. Its estimated costs, by quarter, for the coming year are given below:

   

Quarter
   First Second Third Fourth
Direct materials $ 160,000 $ 80,000 $ 40,000 $ 120,000
Direct labor 160,000 80,000 40,000 120,000
Manufacturing overhead 220,000 196,000 184,000 ?
Total manufacturing costs (a) $ 540,000 $ 356,000 $ 264,000 $ ?
Number of units to be produced (b) 160,000 80,000 40,000 120,000
Estimated unit product cost (a) ÷ (b) $ 3.38 $ 4.45 $ 6.60 $ ?

Management finds the variation in quarterly unit product costs to be confusing and difficult to work with. It has been suggested that the problem lies with manufacturing overhead because it is the largest element of total manufacturing cost. Accordingly, you have been asked to find a more appropriate way of assigning manufacturing overhead cost to units of product.

Required:

1. Assuming the estimated variable manufacturing overhead cost per unit is $0.30, what must be the estimated total fixed manufacturing overhead cost per quarter?

2. Assuming the assumptions about cost behavior from the first three quarters hold constant, what is the estimated unit product cost for the fourth quarter?

3. What is causing the estimated unit product cost to fluctuate from one quarter to the next?

4. Assuming the company computes one predetermined overhead rate for the year rather than computing quarterly overhead rates, calculate the unit product cost for all units produced during the year.
Answer #1

Total fixed manufacturing overhead cost = Total overhead cost – variable overhead cost

Using first quarter, Fixed overhead = 220,000 – 160,000*0.30

= $172,000

2.Manufacturing overhead = 172000+0.3*120,000 = $208000

Unit product cost = (120,000+120,000+208000)/120,000

= $3.73333

i.e. $3.73 per unit

3.Diffence in Number of units produced is leading to a difference in fixed manufacturing overhead per unit and hence, difference in unit product cost

4.Rate = (172000*4)/400,000+0.30

= $2.02 per unit

Unit cost = 1+1+2.02

= $4.02 per unit